OpenStudy (anonymous):

Help! how do you graph 0=a^2+bx+c??? like if you're trying to solve it/find the roots?!?! thats all i really need to know, is how to find the roots. helo helppp

myininaya (myininaya):

bx=-c-a^2 x={-c-a^2}/b gives us the line's x intercept

OpenStudy (anonymous):

Google quadratic formula.

myininaya (myininaya):

its 0=a^2+bx+c this is not a quadratic

OpenStudy (anonymous):

I dont get it, i googled and stuff, all i need to know how to do is find the roots.

OpenStudy (anonymous):

how about this.: 4x^2+x+3=0. how do i find the roots?

myininaya (myininaya):

thats a quadratic

myininaya (myininaya):

a=4 b=1 c=3 if you want you can use the quadratic forumla you can try to factor you can even complete the square to solve for x

myininaya (myininaya):

completing the square for ax^2+bx+c=0 is how the quadratic formula is derived